A vacant city lot is being turned into a neighborhood garden. The neighbors want to fence in a triangular section of the lot and plant flowers there. The longest side of the triangular section is 8 feet shorter than twice the shortest side. The third side is 4 feet longer than the shortest side. The perimeter is 56 feet. How long is each​ side?
Answer by ankor@dixie-net.com(22740)

let s = the length of the shortest side, side 1
:
Write an equation for each statement
:
" The longest side of the triangular section is 8 feet shorter than twice the shortest side."
side 2 = 2s - 8
" The third side is 4 feet longer than the shortest side."
side 3 = s + 4
:
How long is each side?
the sum of the 3 sides is 56 ft
(2s-8) + (s+4) + s = 56
combine like terms
4s - 4 = 56
4s = 56 + 4
4s = 60
s = 60/4
s = 15 ft is the shortest side
:
I'll let you find the other two sides, check that they add up to 56